OpenStudy (anonymous):

A car of mass 800kg moving at 20ms^-1 is brought to rest by application of the brakes in a distance of 100m. Find the force exerted by the brake(friction) if there is no other resistance to motion.

OpenStudy (anonymous):

@Algebraic! @Callisto @physicswork

OpenStudy (anonymous):

Vf^2 =Vi^2 -2ad 0 =Vi^2 -2ad Vi^2 = 2ad a= Vi^2/2d F=ma

OpenStudy (anonymous):

i got the answer thnkss :)